# Generalized Sobolev Transport for Probability Measures on a Graph

### v1

@@ Abstract
We study the optimal transport ($*{}^{*}$: Two authors contributed equally. problem for measures supported on a graph metric space. Recently, Le et al. (2022) leverage the graph structure and propose a variant of OT, namely Sobolev transport (ST), which yields a closed-form expression for a fast computation. However, ST is essentially coupled with the $L^{p}$ geometric structure within its definition which makes it nontrivial to utilize ST for other prior structures. In contrast, the classic OT has the flexibility to adapt to various geometric structures by modifying the underlying cost function. An important instance is the Orlicz-Wasserstein (OW) which moves beyond the $L^{p}$ structure by leveraging the Orlicz geometric structure. Comparing to the usage of standard $p$-order Wasserstein, OW remarkably helps to advance certain machine learning approaches. Nevertheless, OW brings up a new challenge on its computation due to its two-level optimization formulation. In this work, we leverage a specific class of convex functions for Orlicz structure to propose the generalized Sobolev transport (GST). GST encompasses the ST as its special case, and can be utilized for prior structures beyond the $L^{p}$ geometry. In connection with the OW, we show that one only needs to simply solve a univariate optimization problem to compute the GST, unlike the complex two-level optimization problem in OW. We empirically illustrate that GST is several-order faster than the OW. Moreover, we provide preliminary evidences on the advantages of GST for document classification and for several tasks in topological data analysis.
@@ 1 Introduction
Optimal transport (OT) is a natural geometry to compare probability measures (Villani, 2008; Peyré & Cuturi, 2019). Intuitively, OT lifts the cost metric on support data points of input measures to the distance for the measures by minimizing the cost to transport a measure into the other. OT has been applied on many applications in various research fields, e.g., in machine learning (Nadjahi et al., 2019; Titouan et al., 2019; Janati et al., 2020; Mukherjee et al., 2021; Altschuler et al., 2021; Fatras et al., 2021; Scetbon et al., 2021; Le et al., 2021; Takezawa et al., 2022; Fan et al., 2022; Bunne et al., 2022, 2023a; Bonet et al., 2023; Mahey et al., 2023; Nguyen et al., 2023a, b; Korotin et al., 2023, 2024; Nguyen et al., 2024), statistics (Mena & Niles-Weed, 2019; Weed & Berthet, 2019; Liu et al., 2022; Nguyen et al., 2022; Nietert et al., 2022, 2023; Wang et al., 2022), and biology (Schiebinger et al., 2019; Tong et al., 2020; Bunne et al., 2023a, b; Korotin et al., 2024), to name a few.
@@
  • (i) We leverage a certain class of convex functions corresponding to Orlicz geometric structure, and propose the GST for probability measures supported on a graph metric space. We show that GST can be computed by simply solving a univariate optimization problem.
  • (ii) We demonstrate that ST is a special case of the proposed GST. Additionally, GST utilizes the Orlicz geometric structure in the same sense as the OW for OT problem. We further draw a connection between GST and OW.
  • (iii) We empirically illustrate that GST is more computationally efficient than OW. We also show some preliminary evidences on the advantages of GST for document classification and for several tasks in topological data analysis (TDA).
Organization. We briefly review related notations used in the development of our proposals in §2. In §3, we describe our proposed approach, the generalized Sobolev transport (GST). We show that GST is a metric, and then derive its properties and draw its connection to ST and OW in §4. In §5, we empirically illustrate the computational advantages of GST over OW, and show preliminary improvements of GST on document classification and TDA. In §6, we give concluding remarks. Finally, we defer the proofs of key theoretical results and additional materials to the Appendices.
@@ 2.3 Orlicz functional space and Orlicz-Wasserstein
$W_{\Phi}(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\inf\Big{[}t>0:\int_{{\mathbb{G}}\times{\mathbb{G}}}\Phi\left(\frac{d_{{\mathbb{G}}}(x,z)}{t}\right)\text{d}\pi(x,z)\leq 1\Big{]},$ (4)
@@ 3 Generalized Sobolev Transport (GST)
Unlike the OW which directly utilizes the given $N$-function within the Luxemburg norm of the Orlicz functional space (Equation (4)), in order to generalize the ST, we also require a notion of the complementary function of the given $N$-function.
@@
[…]
@@ 4 Properties of the GST
Let $\Phi(t):=\frac{(p-1)^{p-1}}{p^{p}}t^{p}$ with $1<p<\infty$, and let $\omega$ be a nonnegative Borel measure on graph ${\mathbb{G}}$. Then the GST admits the following closed-form expression:
[…]
@@
Connection with Orlicz-Wasserstein and OT. In the special case $\Phi(t)=t^{p}$ with $p\geq 1$, the OW coincides with the $p$-order Wasserstein (Guha et al., 2023). [^9] The $p$-order Wasserstein is reviewed in §B.5. Although $\Phi(t)=t$ is not an $N$-function due to its linear growth, it can be regarded as the limit $p\to 1^{+}$ for the function $\Phi(t)=t^{p}$. In the similar way, Andoni et al. (2018) leverages the Huber function with linear growth for $\Phi(t)$ for Orlicz loss regression in applications. It is worth noting that OW in Equation (4) is a two-level optimization problem involving the transportation plan and the positive scalar $t$. Thus, it is more challenging to compute the OW than the classic OT (e.g., with squared Euclidean cost).
@@
Notice that our obtained results show that one only needs to solve a univariate optimization problem to compute the GST. On the other hand, it is much more challenging to solve the two-level optimization problem (4) for the OW. Notably, a recent realization is that solving exactly OT problems leads to overfitting (Peyré & Cuturi, 2019, §8.4). Therefore, it would be self-defeating to give excessive efforts to optimize the OT problems since it would lead to overfitting within the computation of the OT problems themselves. Consequently, GST can be regarded as a regularization approach for OW.
@@ 5 Experiments
Figure 1: Time consumption for GST and OW on ${\mathbb{G}}_{\text{Log}}$.
Figure 2: Time consumption for GST and OW on ${\mathbb{G}}_{\text{Sqrt}}$.
@@ 5.1 Computation
We compare the time consumption of GST and OW with popular $N$-functions $\Phi_{1},\Phi_{2}$, and the limit case $\Phi_{0}$.
@@
Figure 4: Document classification on graph ${\mathbb{G}}_{\text{Sqrt}}$.
@@ 6 Conclusion
We propose the generalized Sobolev transport (GST) for probability measures supported on a graph metric space. This is achieved by leveraging a special family of convex functions (i.e., the set of $N$-functions) to extend the Sobolev transport (ST), which is a scalable variant of OT on a graph. This novel approach enables us to adopt Orlicz geometric structure for GST. Consequently, GST can inherit advantages from both ST and the Orlicz-Wasserstein (OW). An important feature is that unlike OW which involves a complex two-level optimization problem, one can simply solve a univariate optimization problem for the GST computation.
@@ B.4 Further discussions
Orlicz geometric structure. For regression problem, squared $\ell_{2}$ norm is the classic loss, i.e., the least squared regression ($\ell_{2}$ regression). An alternative choice for the loss function is $\ell_{1}$ norm, i.e., least absolute deviation regression ($\ell_{1}$ regression) which gives more robust solution (Gorard, 2005). A more generalized class of loss functions stems from $\mathbb{M}$-estimator, i.e., $l_{\mathbb{M}}(x)=\sum_{i}\mathbb{M}(x_{i})$ where $\mathbb{M}$ is an $\mathbb{M}$-estimator function (e.g., Huber, Cauchy, Welsh, Tukey, Geman-McClure functions (Zhang, 1997, Table 1)). When $\mathbb{M}$-estimator is the Huber function, it is similar to $\ell_{2}$ regression when $t$ is small, and looks like $\ell_{1}$ regression when $t$ is large. Therefore, it can enjoy both benefits from the $\ell_{1}$ and $\ell_{2}$ regressions. However, such $\mathbb{M}$-estimation functional loss depends on the data scale, i.e., one may get different solutions if the data are rescaled. Such issue can be addressed by using the normalized $\mathbb{M}$-estimation function loss (Andoni et al., 2018), i.e., a scale-free version of $\mathbb{M}$-estimator.
Deng et al. (2022) propose an efficient data structure for the Distance Oracle problem including Orlicz norm as a special case which helps to scale up many machine learning problems, e.g., reinforcement learning, kernelized support vector machine, and clustering.
Chamakh et al. (2020) leverage $\alpha$-exponential Orlicz on the sample distribution to derive a finite-sample deviation bound for a general class of polynomial-growth functions. Consequently, one can address the challenging polynomial growing nature of the underlying function class in the empirical process for random Fourier features on approximation of high-order kernel derivatives. In particular, Chamakh et al. (2020) propose finite-sample uniform guarantee for random Fourier features which is almost surely convergence to approximate high-order derivatives for arbitrary kernel.
@@ B.5 Further empirical results
Figure 11: Topological data analysis $M=10^{3}$ on graph ${\mathbb{G}}_{\text{Log}}$.
Figure 12: Topological data analysis $M=10^{3}$ on graph ${\mathbb{G}}_{\text{Sqrt}}$.
Figure 13: Topological data analysis $M=10^{2}$ on graph ${\mathbb{G}}_{\text{Log}}$.
Figure 14: Topological data analysis $M=10^{2}$ on graph ${\mathbb{G}}_{\text{Sqrt}}$.

### v2

@@ Abstract
We study the optimal transport (OT) problem for measures supported on a graph metric space. Recently, Le et al. (2022) leverage the graph structure and propose a variant of OT, namely Sobolev transport (ST), which yields a closed-form expression for a fast computation. However, ST is essentially coupled with the $L^{p}$ geometric structure within its definition which makes it nontrivial to utilize ST for other prior structures. In contrast, the classic OT has the flexibility to adapt to various geometric structures by modifying the underlying cost function. An important instance is the Orlicz-Wasserstein (OW) which moves beyond the $L^{p}$ structure by leveraging the Orlicz geometric structure. Comparing to the usage of standard $p$-order Wasserstein, OW remarkably helps to advance certain machine learning approaches. Nevertheless, OW brings up a new challenge on its computation due to its two-level optimization formulation. In this work, we leverage a specific class of convex functions for Orlicz structure to propose the generalized Sobolev transport (GST). GST encompasses the ST as its special case, and can be utilized for prior structures beyond the $L^{p}$ geometry. In connection with the OW, we show that one only needs to simply solve a univariate optimization problem to compute the GST, unlike the complex two-level optimization problem in OW. We empirically illustrate that GST is several-order faster than the OW. Moreover, we provide preliminary evidences on the advantages of GST for document classification and for several tasks in topological data analysis.
Scalability, Generalized Sobolev Transport, Measures on a Graph, Lp geometric structure, Orlicz geometric structure, Sobolev Transport, Orlicz-Wasserstein
@@ 1 Introduction
Optimal transport (OT) is a natural geometry to compare probability measures (Villani, 2008; Peyré & Cuturi, 2019). Intuitively, OT lifts the cost metric on support data points of input measures to the distance for the measures by minimizing the cost to transport a measure into the other. OT has been applied on many applications in various research fields, e.g., in machine learning (Nadjahi et al., 2019; Titouan et al., 2019; Janati et al., 2020; Mukherjee et al., 2021; Altschuler et al., 2021; Fatras et al., 2021; Scetbon et al., 2021; Le & Nguyen, 2021; Le et al., 2021a, b; Liu et al., 2021; Nguyen et al., 2021a, b; Takezawa et al., 2022; Fan et al., 2022; Bunne et al., 2022, 2023a; Bonet et al., 2023; Mahey et al., 2023; Nguyen et al., 2023a, b; Hua et al., 2023; Korotin et al., 2023, 2024; Le et al., 2024; Nguyen et al., 2024), statistics (Mena & Niles-Weed, 2019; Weed & Berthet, 2019; Liu et al., 2022; Nguyen et al., 2022; Wang et al., 2022; Nietert et al., 2022, 2023; Pham et al., 2024), and biology (Schiebinger et al., 2019; Tong et al., 2020; Bunne et al., 2023a, b; Korotin et al., 2024).
@@
  • We leverage a certain class of convex functions corresponding to Orlicz geometric structure, and propose the GST for probability measures supported on a graph metric space. We show that GST can be computed by simply solving a univariate optimization problem.
  • We demonstrate that ST is a special case of the proposed GST. Additionally, GST utilizes the Orlicz geometric structure in the same sense as the OW for OT problem. We further draw a connection between GST and OW.
  • We empirically illustrate that GST is more computationally efficient than OW. We also show some preliminary evidences on the advantages of GST for document classification and for several tasks in topological data analysis (TDA).
Organization. We briefly review related notations used in the development of our proposals in §2. In §3, we describe our proposed approach, the generalized Sobolev transport (GST). We show that GST is a metric, and then derive its properties and draw its connection to ST and OW in §4. In §5, we empirically illustrate the computational advantages of GST over OW, and show preliminary improvements of GST on document classification and TDA. In §6, we give concluding remarks. Additionally, we defer the proofs of key theoretical results and additional materials to the Appendices. We have also released code for our proposed approach. [^2] The code repository is on https://github.com/lttam/Generalized-Sobolev-Transport
@@ 2.3 Orlicz functional space and Orlicz-Wasserstein
[…]
$\int_{{\mathbb{G}}\times{\mathbb{G}}}\Phi\left(\frac{d_{{\mathbb{G}}}(x,z)}{t}\right)\text{d}\pi(x,z)\leq 1\Big{]},$ (4)
@@ 3 Generalized Sobolev Transport (GST)
Unlike the OW which directly utilizes the given $N$-function within the Luxemburg norm of the Orlicz functional space (Equation (2.3)), in order to generalize the ST, we also require a notion of the complementary function of the given $N$-function.
@@
Let $\text{supp}(\mu)$ be the set of supports of $\mu$. Observe that for any support $x\in\text{supp}(\mu)$, its mass is accumulated into $\mu(\gamma_{e})$ if and only if $e\subset[z_{0},x]$. Let define set $E_{\mu,\nu}\subset E$ as
$E_{\mu,\nu}:=\left\{e\!\in\!E\mid\exists z\!\in\!(\text{supp}(\mu)\cup\text{supp}(\nu)),e\subset[z_{0},z]\right\}\!.$
@@ 4 Properties of the GST
Let $\Phi(t):=\frac{(p-1)^{p-1}}{p^{p}}t^{p}$ with $1<p<\infty$, and let $\omega$ be a nonnegative Borel measure on graph ${\mathbb{G}}$. Then, the GST admits the following closed-form expression:
$\mathcal{GS}_{\Phi}(\mu,\nu)=\left(\int_{{\mathbb{G}}}\left|\mu(\Lambda(x))-\nu(\Lambda(x))\right|^{p}\omega(\text{d}x)\right)^{\frac{1}{p}},$ (9)
@@
Connection with Orlicz-Wasserstein and OT. In the special case $\Phi(t)=t^{p}$ with $p\geq 1$, the OW coincides with the $p$-order Wasserstein (Guha et al., 2023). [^10] The $p$-order Wasserstein is reviewed in §B.5. Although $\Phi(t)=t$ is not an $N$-function due to its linear growth, it can be regarded as the limit $p\to 1^{+}$ for the function $\Phi(t)=t^{p}$. In the similar way, Andoni et al. (2018) leverages the Huber function with linear growth for $\Phi(t)$ for Orlicz loss regression in applications. It is worth noting that OW in Equation (2.3) is a two-level optimization problem involving the transportation plan and the positive scalar $t$. Thus, it is more challenging to compute the OW than the classic OT (e.g., with squared Euclidean cost).
@@
Notice that our obtained results show that one only needs to solve a univariate optimization problem to compute the GST. On the other hand, it is much more challenging to solve the two-level optimization problem (2.3) for the OW. Notably, a recent realization is that solving exactly OT problems leads to overfitting (Peyré & Cuturi, 2019, §8.4). Therefore, it would be self-defeating to give excessive efforts to optimize the OT problems since it would lead to overfitting within the computation of the OT problems themselves. Consequently, GST can be regarded as a regularization approach for OW.
@@ 5 Experiments
Figure 1: Time consumption for GST and OW on ${\mathbb{G}}_{\text{Log}}$.
Figure 2: Time consumption for GST and OW on ${\mathbb{G}}_{\text{Sqrt}}$.
@@
Figure 4: Document classification on graph ${\mathbb{G}}_{\text{Sqrt}}$.
@@ 5.1 Computation
We compare the time consumption of GST and OW with popular $N$-functions $\Phi_{1},\Phi_{2}$, and the limit case $\Phi_{0}$. [^14] Notice that GST-$\Phi_{0}$ is equal to $1$-order Sobolev transport (see Remark 4.5), and OW-$\Phi_{0}$ is equal to the OT on a graph (i.e., OT with graph metric ground cost).
@@ 6 Conclusion
We propose the generalized Sobolev transport (GST) for probability measures supported on a graph metric space. This is achieved by leveraging a special family of convex functions (i.e., the set of $N$-functions) to extend the Sobolev transport (ST), which is a scalable variant of OT on a graph. This novel approach enables us to adopt Orlicz geometric structure for GST. Consequently, GST can inherit advantages from both ST and the Orlicz-Wasserstein (OW). An important feature is that unlike OW which involves a complex two-level optimization problem, one can simply solve a univariate optimization problem for the GST computation. Thus, GST is a computationally efficient variant for the computationally-demanding OW, which paves a way to use OW in applications. For future work, it is interesting to extend the GST for measures possibly having different total mass (i.e., unbalanced setting), which is essential for applications having noisy supports or outliers in input measures (Frogner et al., 2015; Balaji et al., 2020; Mukherjee et al., 2021; Fatras et al., 2021).
Acknowledgements
We thank the area chairs and anonymous reviewers for their comments. KF has been supported in part by Grant-in-Aid for Transformative Research Areas (A) 22H05106. TL gratefully acknowledges the support of JSPS KAKENHI Grant number 23K11243, and Mitsui Knowledge Industry Co., Ltd. grant.
Impact Statement
The paper presents a generalized version for the Sobolev transport (ST) which is a scalable variant of optimal transport for probability measures on a graph. By leveraging a special family of convex functions, we can adopt Orlicz geometric structure for the generalized Sobolev transport (GST). Furthermore, GST can be regarded as a computationally efficient variant of the computationally-demanding Orlicz-Wasserstein (OW), which possesses implicit essences to advance certain results in machine learning. Therefore, the proposed GST can inherit the merits from both ST and OW. To our knowledge, there are no foresee potential societal consequences of our work.
@@ Appendix A Detailed Proofs
$\mathcal{GS}_{\Phi}(\mu,\nu)=\inf_{k>0}\frac{1}{k}\left(1+\int_{{\mathbb{G}}}\Phi(k\left|h(x)\right|)\omega(\text{d}x)\right)$ (16)
@@ B.4 Further discussions
OT problems for measures on a graph. In this work, our considered problem is to compute OT distance between two input probability measures supported on the same graph. We distinguish this line of research to the OT problem between two (different) input graphs, considered in Petric Maretic et al. (2019); Dong & Sawin (2020); Tam & Dunson (2022), where its goal is to compute the distance between two input graphs by using OT approach.
@@ B.5 Further empirical results
Figure 11: Topological data analysis on graph ${\mathbb{G}}_{\text{Log}}$ ($M=10^{3}$).
Figure 12: Topological data analysis on graph ${\mathbb{G}}_{\text{Sqrt}}$ ($M=10^{3}$).
Figure 13: Topological data analysis on graph ${\mathbb{G}}_{\text{Log}}$ ($M=10^{2}$).
Figure 14: Topological data analysis on graph ${\mathbb{G}}_{\text{Sqrt}}$ ($M=10^{2}$).